# New Projection Operators for Planar Electrodynamics

### v1

@@ 6.2   Analysis for the Maxwell–Lee–Wick–Chern–Simons model
$p^{2}=0,\qquad p^{2}\!\left(1-\frac{p^{2}}{m^{2}}\right)^{2}-\kappa^{2}=0.$ (44)
The pole at $p^{2}=0\Rightarrow p^{0}=\pm\|\vec{p}\|$ is purely a gauge pole and disappears when the propagator is saturated with a conserved current, so that the physical structure is determined by the cubic term in $p^{2}$. For each fixed $\vec{p}$, defining $x=((p^{0})^{2}-\vec{p}^{\,2})/m^{2}$ and $\beta\equiv\kappa^{2}/m^{2}$,  (44) becomes
@@
  • $\frac{\kappa^{2}}{m^{2}}<\frac{4}{27}$: three real massive modes (one of them with negative norm, as we will see below);
  • $\frac{\kappa^{2}}{m^{2}}=\frac{4}{27}$: critical case, with two degenerate modes;
  • $\frac{\kappa^{2}}{m^{2}}>\frac{4}{27}$: one real massive mode and one pair of complex-conjugate poles (Lee–Wick pair).
Each real root $x_{i}$ define $p_{i}^{2}=m^{2}x_{i}$. Therefore the dispersion relationship is $p^{0}=\pm\sqrt{\vec{p}^{\,2}+p_{i}^{2}}$. In other words, the physical modes are relativistic.
@@ 6.2.1  Saturated amplitude and residues
For a conserved current, the longitudinal sector ($P_{1}$) does not contribute. Using  (25) with the coefficients above, the saturated amplitude is
$A(p)=\frac{\|J^{(+)}\|^{2}}{-(p^{0})^{2}+\vec{p}^{\,2}+\frac{\big((p^{0})^{2}-\vec{p}^{\,2}\big)^{2}}{m^{2}}-\kappa\,\sqrt{(p^{0})^{2}-\vec{p}^{\,2}}}$
$\frac{\|J^{(-)}\|^{2}}{-(p^{0})^{2}+\vec{p}^{\,2}+\frac{\big((p^{0})^{2}-\vec{p}^{\,2}\big)^{2}}{m^{2}}+\kappa\,\sqrt{(p^{0})^{2}-\vec{p}^{\,2}}}.$ (46)
@@
Defining the product of denominators $D(p^{0},\vec{p})=\left[(p^{0})^{2}-\vec{p}^{\,2}\right]G(p^{2})$, the residue of the saturated amplitude at $p^{0}=E^{*}$ is
$\text{Res}\,A\Big|_{p^{2}=\lambda}=-\,\frac{1-\lambda/m^{2}}{2G^{\prime}(\lambda)E^{*}}.$ (47)
Since,
@@
$G^{\prime}(\lambda)=\left(1-\frac{\lambda}{m^{2}}\right)\left(1-\frac{3\lambda}{m^{2}}\right).$ (48)
@@
How $E^{*}>0$, we see that the sign of the residue is controlled by
@@
$\text{sgn}\,\bigl(\text{Res}\,A\bigr)=\text{sgn}\!\left[-\,\frac{1}{1-3\lambda/m^{2}}\right].$ (49)
Thus, for modes with $0<\lambda<m^{2}/3$ the residue is negative, characterizing a ghost mode (Lee–Wick partner), whereas modes with $\lambda>m^{2}/3$ have positive residue and correspond to physical unitary degrees of freedom. Adopting an appropriate contour prescription in the complex $p^{0}$ plane, these modes can be handled in such a way that the SS–matrix remains unitary and macroscopic causality is preserved, while possible microscopic violations of causality are confined to distances of order $1/\Gamma_{n}$ and are suppressed by the high mass scales characterizing the Lee–Wick partners.
@@ 6.2.2  Causality
For parameter values such that all poles $\lambda$ are real and positive ($\kappa^{2}/m^{2}<4/27$), the associated modes have relativistic dispersion relations and subluminal group velocities, preserving microscopic causality. When $\kappa^{2}/m^{2}>4/27$, two of the poles become complex conjugates; in this situation the coordinate-space propagator exhibits damped oscillations, corresponding to unstable resonances. Following the Lee–Wick contour prescription, one chooses the integration contour in $p^{0}$ so that such modes decay in the future, preserving the unitarity of the $S$-matrix and macroscopic causality, at the cost of small local violations of microcausality, suppressed by the Lee–Wick mass scale.
@@ 6.3   Analysis for the Maxwell–Deser–Jackiw model
which describes a single massive mode with mass $m>0$ and relativistic energy $E(\vec{p})$.
@@ 6.3.1  Saturated amplitude and residues
The saturated amplitude with a conserved current is obtained with $J=\theta J$ and $J^{*}\!\cdot\omega J=0$:
$A(p)=-\,\frac{1}{p^{2}\!\left(1-\dfrac{p^{2}}{m^{2}}\right)}\left(J^{*}\!\cdot J+\frac{i}{m}\,J^{*}\!\cdot S\cdot J\right).$ (52)
In the transverse subspace, we diagonalize $S$ in the basis
$SJ^{(\pm)}=\pm i\sqrt{p^{2}}\,J^{(\pm)},\qquad J=J^{(+)}+J^{(-)}.$ (53)
Hence,
$J^{*}\!\cdot J$ $=\|J^{(+)}\|^{2}+\|J^{(-)}\|^{2},$ (54)
$J^{*}\!\cdot S\cdot J$ $=+i\sqrt{p^{2}}\,\|J^{(+)}\|^{2}-i\sqrt{p^{2}}\,\|J^{(-)}\|^{2}.$
Substituting into (52),
$A(p^{0},\vec{p})=-\frac{1}{d(p^{0},\vec{p})}\Bigg[\|J^{(+)}\|^{2}\left(1-\frac{\sqrt{(p^{0})^{2}-\vec{p}^{\,2}}}{m}\right)\Bigg]$
[…]
At the physical pole we have
$A(p^{0},\vec{p})\Big|_{(p^{0})^{2}\simeq\vec{p}^{\,2}+m^{2}}\simeq-\,\frac{1}{d(p^{0},\vec{p})}\Big[\,0\cdot\|J^{(+)}\|^{2}+2\|J^{(-)}\|^{2}\Big].$ (56)
Where
$d(p^{0},\vec{p})=\big((p^{0})^{2}-\vec{p}^{\,2}\big)\left(1-\dfrac{(p^{0})^{2}-\vec{p}^{\,2}}{m^{2}}\right).$
Furthermore, one can also write
$A(p^{0},\vec{p})\Big|_{(p^{0})^{2}\simeq\vec{p}^{\,2}+m^{2}}\simeq\frac{\|J^{(-)}\|^{2}}{E^{*}(\vec{p})}\,\frac{1}{p^{0}-E^{*}(\vec{p})}$ (57)
This shows that only the negative term propagates, whereas the positive is projected out on the mass shell. The residue at $p^{o}=E^{*}(\vec{p})=\sqrt{\vec{p}^{\,2}+m^{2}}$ is
$\operatorname*{Res}_{\,p^{0}=E^{*}(\vec{p})}A(p^{0},\vec{p})=\frac{\|J^{(-)}\|^{2}}{E^{*}(\vec{p})}.$ (58)
How $E^{*}>0$ this demonstrates that the massive mode has positive norm and that the model is unitary at tree level.
@@ 6.3.2  Causality
$(p^{0})^{2}-\vec{p}^{\,2}=m^{2}\Rightarrow p^{0}=\pm E(\vec{p})\Rightarrow E(\vec{p})=\sqrt{\vec{p}^{\,2}+m^{2}}.$ (59)
@@
$v_{g}=\left|\frac{\partial E(\vec{p})}{\partial|\vec{p}|}\right|=\frac{|\vec{p}|}{\sqrt{\vec{p}^{\,2}+m^{2}}}<1,$ (60)
@@ Appendix A Supplementary Material
Therefore, we see that the vectors in $\text{Im}\,Z$ and $\ker Z$ are eigenvectors of $S$. Now, let us show that $\text{Im}\,Z$ and $\ker Z$ cannot be trivial subspaces! For this to be true, we must have $\dim(\text{Im}\,Z)=\dim(\ker Z)=1$. Suppose initially that one of them is trivial and we will arrive at a contradiction.
Case 1: $\text{Im}\,Z=W$. Thus, $\forall\,w\in W\implies\text{Im}\,Z=W$, i.e.
@@
Case 2: $\ker Z=W$. Thus, $\forall\,w\in W\implies\ker P_{W}=W$, i.e.
@@
In general, $S_{W}=a\mathbf{1}_{W}$. If $v=\omega(v)+\theta(v)$, then
@@
$S(v)=a\mathbf{1}_{W}(\theta(v))\implies S(v)=a\theta(v),$
@@
Since $\theta$ is symmetric, we can write
$S_{\mu\nu}=a\theta_{\nu\mu},$
but
$S_{\mu\nu}=S_{\nu\mu},\quad S_{\mu\nu}=-S_{\nu\mu},$
which implies $S_{\mu\nu}=0$, that is, $\theta=0$. But this is absurd, since $\theta\neq 0$. $\blacksquare$

### v2

@@ 6.2  Analysis for the Maxwell–Lee–Wick–Chern–Simons model
$p^{2}=0,\qquad p^{2}\!\left(1-\mu p^{2}\right)^{2}-\kappa^{2}=0.$ (46)
The pole at $p^{2}=0\Rightarrow p^{0}=\pm\|\vec{p}\|$ is purely a gauge pole and disappears when the propagator is saturated with a conserved current, so that the physical structure is determined by the cubic term in $p^{2}$. In this sense, it is important to note that at $p^{2}=0$ the standard decompositions based on $\omega_{\mu\nu}$ and on the projector basis associated with the transverse subspace are no longer uniformly well defined. Consequently, the analysis of the massless pole must be carried out either by a careful limiting procedure or by adopting a light-like parametrization of the external sources. Although the propagator exhibits explicit $1/p^{2}$ factors, including in terms proportional to $S_{\mu\nu}$, saturation with conserved currents removes the pure gauge sector ($\omega_{\mu\nu}$), while the contributions involving $S_{\mu\nu}$ are suppressed, since $S_{\mu\nu}\sim\epsilon_{\mu\nu\rho}p^{\rho}$. The remaining $1/p^{2}$ behavior in the saturated amplitude $\mathcal{A}(p)$ should therefore be interpreted as a long-range interaction rather than as a genuine propagating massless degree of freedom, and the appropriate criterion for massless unitarity is the evaluation of the sources together with the gauge-induced constraints, as in the approach of [1]. For each fixed $\vec{p}$, defining $x=\mu((p^{0})^{2}-\vec{p}^{\,2})$ and $\beta\equiv\kappa^{2}\mu$,  (46) becomes
@@
  • $\kappa^{2}\mu<\frac{4}{27}$: three real massive modes (two of them with negative residue, as we will see below);
  • $\kappa^{2}\mu=\frac{4}{27}$: critical case, with two degenerate modes;
  • $\kappa^{2}\mu>\frac{4}{27}$: one real massive mode and one pair of complex-conjugate poles (Lee–Wick pair);
  • $\kappa^{2}\mu<0$: a single real massive mode in the physical region $p^{2}>0$.
Each real root $x_{i}$ define $p_{i}^{2}=x_{i}/\mu$. Therefore the dispersion relationship is $p^{0}=\pm\sqrt{\vec{p}^{\,2}+p_{i}^{2}}$. In other words, the physical modes correspond to relativistic particles with well-defined masses $m_{i}=\sqrt{x_{i}/\mu}$.
@@ 6.2.1  Saturated amplitude and residues
For a conserved current, the longitudinal sector ($P_{1}$) does not contribute to the amplitude. Using  (25) with the coefficients above, the saturated amplitude is
$\mathcal{A}(p)=\frac{\|J^{(+)}\|^{2}}{(p^{0})^{2}-\vec{p}^{\,2}-\mu\big((p^{0})^{2}-\vec{p}^{\,2}\big)^{2}-\kappa\,\sqrt{(p^{0})^{2}-\vec{p}^{\,2}}}$
$\frac{\|J^{(-)}\|^{2}}{(p^{0})^{2}-\vec{p}^{\,2}-\mu\big((p^{0})^{2}-\vec{p}^{\,2}\big)^{2}+\kappa\,\sqrt{(p^{0})^{2}-\vec{p}^{\,2}}}.$ (48)
@@
Defining the product of denominators $D(p^{0},\vec{p})=\left[(p^{0})^{2}-\vec{p}^{\,2}\right]G(p^{2})$, the residue of the saturated amplitude at $p^{0}=E_{*}(\vec{p})$ is
@@
$\operatorname*{Res}_{\,p^{0}=E_{*}(\vec{p})}\mathcal{A}(p^{0},\vec{p})=\frac{\left(1-\mu\lambda\right)}{G^{\prime}(\lambda)E_{*}(\vec{p})}\|J^{(\sigma)}\|^{2},\qquad\sigma=\begin{cases}+,&\kappa\left(1-\mu\lambda\right)>0,\\ -,&\kappa\left(1-\mu\lambda\right)<0.\end{cases}$ (49)
@@
Where $G^{\prime}(\lambda)=\left(1-\mu\lambda\right)\left(1-3\mu\lambda\right).$ Since $E_{*}(\vec{p})>0$, we see that the sign of the residue is controlled by
@@
$\text{sgn}\,\bigl(\text{Res}\,\mathcal{A}\bigr)=\text{sgn}\!\left[\,\frac{1}{1-3\mu\lambda}\right].$ (50)
Now, consider $D_{\pm}(p)=-p^{2}(1-\mu p^{2})\pm\kappa\sqrt{p^{2}},$ in the physical region $p^{2}>0$. Introducing $t=\sqrt{p^{2}}>0$, we rewrite
$D_{\pm}(p)=-t^{2}(1-\mu t^{2})\pm\kappa t=t\bigl[\mu t^{3}-t\pm\kappa\bigr].$ (51)
Hence, besides the massless factor $t=0$, the massive poles in $p^{2}>0$ are determined by the cubic equations
$f_{\pm}(t):=\mu t^{3}-t\pm\kappa=0,\qquad t>0,$ (52)
whose dependence on the parameters $\mu$ and $\kappa$ is analysed below. Considering the auxiliary function $f_{0}(t)=t\,(\mu\,t^{2}-1)$, and the result $f^{\prime}_{\pm}(t)=f^{\prime}_{0}(t)=3\mu t^{2}-1,$ we obtain the following results that for the corresponding negative $\kappa$ values can be obtained interchanging $f_{+}\leftrightarrow f_{-}$. The Figure 1 summarizes all the results obtained for the pole analysis.
  • In the case $\mu>0$:
    1. 1.
      $0<\kappa<\frac{2}{3\sqrt{3}}\frac{1}{\sqrt{\mu}}$: $f_{+}$ has two positive roots $m_{1}$ and $m_{2}$, whereas $f_{-}$ has only one root $m_{3}$, satisfying the inequalities $\kappa<m_{1}<\frac{1}{\sqrt{3\mu}}<m_{2}<\frac{1}{\sqrt{\mu}}<m_{3}<\frac{2}{\sqrt{3\mu}}$.
    2. 2.
      $\kappa=\frac{2}{3\sqrt{3}}\frac{1}{\sqrt{\mu}}$: $f_{+}$ has a double root $m_{1}=m_{2}=\frac{1}{\sqrt{3\mu}}$, and $f_{-}$ has one root $m_{3}=\frac{2}{\sqrt{3\mu}}$.
    3. 3.
      $\kappa>\frac{2}{3\sqrt{3}}\frac{1}{\sqrt{\mu}}$: $f_{+}$ has no positive root, whereas $f_{-}$ has only one positive root $m_{3}>\frac{2}{\sqrt{3\mu}}$.
  • In the case $\mu\leq 0$:
    1. 1.
      $\kappa>0$: $f_{+}$ has one positive root $m_{*}\leq\kappa$ and $f_{-}$ has no positive root.
[…]
(a) Case $\mu>0$, with $0<\kappa<\frac{2}{3\sqrt{3}}\frac{1}{\sqrt{\mu}}$.
[…]
(b) Case $\mu<0$, with $\kappa>0$.
Figure 1: Representative graphs of the cubic functions $f_{\pm}(t)=\mu t^{3}-t\pm\kappa$ and $f_{0}(t)=t(\mu t^{2}-1)$. (a): where $f_{+}$ has two positive real roots $(m_{1},m_{2})$ and $f_{-}$ has one $(m_{3})$. (b): where a single positive real root is present.
In the three-real-pole regime $0<\kappa^{2}\mu<\frac{4}{27},$ the roots satisfy
$1-\mu m_{1}^{2}>0,\qquad 1-\mu m_{2}^{2}<0,\qquad 1-\mu m_{3}^{2}<0.$ (53)
At a massive pole $p^{2}=m^{2}$ we define $\sigma=\operatorname{sgn}\!\bigl[\kappa\,(1-\mu m^{2})\bigr]$ so that $\sigma=+$ means the pole belongs to $D_{+}(p)=0$ and is saturated by $\|J^{(+)}\|^{2}$, whereas $\sigma=$- means it belongs to $D_{-}(p)=0$ and is saturated by $\|J^{(-)}\|^{2}$. In the three-real-pole regime with (53), one gets
$\sigma_{1}=\sigma_{2}=\operatorname{sgn}(\kappa),\qquad\sigma_{3}=-\operatorname{sgn}(\kappa).$ (54)
For $\mu<0$ the unique massive pole always satisfies $1-\mu m_{*}^{2}>0$, hence
$\sigma_{*}=\operatorname{sgn}(\kappa).$ (55)
The sign of the residue at $p^{2}=m^{2}$ is controlled by $\operatorname{sgn}\!\left(1-3\mu m^{2}\right).$ Thus, in the three-pole regime one finds
$\operatorname{sgn}\!\bigl(\operatorname{Res}\mathcal{A}\bigr)\Big|_{m_{1}}=+,\qquad\operatorname{sgn}\!\bigl(\operatorname{Res}\mathcal{A}\bigr)\Big|_{m_{2},m_{3}}=-,$ (56)
whereas for $\mu<0$ the unique massive pole has $1-3\mu m_{*}^{2}>0$ and therefore positive residue. The derivation of (49) assumes that the physical pole is simple, i.e. $\partial_{0}D(p_{\ast})\neq 0$, or equivalently $G^{\prime}(\lambda_{\ast})\neq 0$. At the critical point $\kappa^{2}\mu=4/27$ the massive root becomes degenerate, satisfying simultaneously
$D_{\sigma}(p_{\ast})=0,\qquad\partial_{0}D_{\sigma}(p_{\ast})=0,$ (57)
which corresponds to a double zero of $G(\lambda)$ at $\lambda_{\ast}=p_{\ast}^{2}=1/(3\mu)$. As a consequence, the linear approximation used in (49) is no longer valid and the transverse propagator develops a second–order pole in $p^{0}$. Indeed, as shown in [1], double poles in the propagator rendering the theory non–unitary in this regime. For completeness, table 2 summarize the pole structure and the corresponding residue signs obtained in the present analysis.
Table 2: Poles structure in the physical region $p^{2}>0$.
Condition Poles in $p^{2}>0$ Sector (vanishing $D_{\pm}$) $\operatorname{sgn}(\mathrm{Res}\,\mathcal{A})$
$\mu>0$
$0<\kappa^{2}\mu<\dfrac{4}{27}$ $m_{1},m_{2},m_{3}$ $\begin{array}[]{l}m_{1,2}:\ \sigma=\operatorname{sgn}(\kappa)\\ m_{3}:\ \sigma=-\operatorname{sgn}(\kappa)\end{array}$ $\begin{array}[]{l}m_{1}:\ +\ (1-3\mu m_{1}^{2}>0)\\ m_{2}:\ -\ (1-3\mu m_{2}^{2}<0)\\ m_{3}:\ -\ (1-3\mu m_{3}^{2}<0)\end{array}$
$\kappa^{2}\mu=\dfrac{4}{27}$ $m_{1}=m_{2}=\dfrac{1}{\sqrt{3\mu}},\ m_{3}$ same criterion as above $m_{1,2}:\text{undefined}\ ;\qquad m_{3}:$-
$\kappa^{2}\mu>\dfrac{4}{27}$ $m_{3}$ $\sigma=-\operatorname{sgn}(\kappa)$ -
$\mu<0$
$\kappa^{2}\mu<0$ $m_{*}$ $\sigma=\,\operatorname{sgn}(\kappa)$ ++
Thus, although a given pole of the transverse propagator may satisfy the positivity condition of the residue in a restricted region of the parameter space, tree-level unitarity requires all propagating modes to have positive residues simultaneously. A detailed analysis of the dispersion relation shows that this condition cannot be fulfilled in the Maxwell–Lee–Wick–Chern–Simons model.
Restricting ourselves to the $\mu>0$ sector, whenever the spectrum contains three real poles, two of them necessarily lie in the region $\lambda>1/3\mu$ and carries a negative residue. Likewise, in the $\mu>0$ regime with a single real pole, the latter is also located at $\lambda>1/3\mu$, again corresponding to a ghost excitation. The sign of $\kappa$ does not affect this conclusion, as it merely selects which transverse contribution saturates the pole, without altering the location of the roots or the sign structure of the residues. Therefore, no choice of parameters renders all propagating modes unitary, and the Maxwell–Lee–Wick–Chern–Simons model is intrinsically nonunitary at tree level.
Finally, let us comment on the $\mu<0$ sector. In this case, the spectrum also contains complex poles, similarly to the $\kappa^{2}\mu>4/27$ regime discussed above. However, the massive real pole that survives in the physical region $p^{2}>0$ carries a positive residue, independently of the sign of $\kappa$. This feature distinguishes the $\mu<0$ sector from the corresponding $\mu>0$ configurations with a single real pole, where the latter is always associated with a ghost excitation. Nevertheless, the presence of complex poles and the lack of a parameter choice yielding a fully healthy spectrum prevent the restoration of tree-level unitarity in the model.
@@ 6.2.2  Causality
For parameter values such that all poles $\lambda$ are real and positive ($\kappa^{2}\mu<4/27$), the associated modes have relativistic dispersion relations and subluminal group velocities, preserving microscopic causality. When $\kappa^{2}\mu>4/27$, two of the poles become complex conjugates; in this situation the coordinate-space propagator exhibits damped oscillations, corresponding to unstable resonances. Following the Lee–Wick contour prescription, one chooses the integration contour in $p^{0}$ so that such modes decay in the future, preserving the unitarity of the $S$-matrix and macroscopic causality, at the cost of small local violations of microcausality, suppressed by the Lee–Wick mass scale.
@@ 6.3  Analysis for the Maxwell–Deser–Jackiw model
which describes a single massive mode with mass $m>0$ and relativistic energy $E_{*}(\vec{p})$.
@@ 6.3.1  Saturated amplitude and residues
Using the projector decomposition introduced in (25), the saturated amplitude can be written as
$\mathcal{A}(p^{0},\vec{p})=\frac{1}{\big((p^{0})^{2}-\vec{p}^{\,2}\big)}\left[\frac{\|J^{(+)}\|^{2}}{\left(1-\frac{\sqrt{(p^{0})^{2}-\vec{p}^{\,2}}}{m}\right)}+\frac{\|J^{(-)}\|^{2}}{\left(1+\frac{\sqrt{(p^{0})^{2}-\vec{p}^{\,2}}}{m}\right)}\right].$ (60)
Near the pole we have
$\mathcal{A}(p^{0},\vec{p})\Big|_{\,p^{0}=E_{*}(\vec{p})}\simeq-\frac{\|J^{(+)}\|^{2}}{E_{*}(\vec{p})}\,\frac{1}{p^{0}-E_{*}(\vec{p})}+\frac{\|J^{(-)}\|^{2}}{2m^{2}}.$ (61)
This shows that only the positive term propagates, whereas the negative is projected out on the mass shell. The residue at $p^{o}=E_{*}(\vec{p})=\sqrt{\vec{p}^{\,2}+m^{2}}$ is
$\operatorname*{Res}_{\,p^{0}=E_{*}(\vec{p})}\mathcal{A}(p^{0},\vec{p})=-\frac{\|J^{(+)}\|^{2}}{E_{*}(\vec{p})}.$ (62)
Since $E_{*}(\vec{p})>0$ and $\|J^{(+)}\|^{2}\geq 0$, the residue is negative. Therefore, the massive mode associated with this pole is a ghost and the model is not unitary at the tree level.
@@ 6.3.2  Causality
$(p^{0})^{2}-\vec{p}^{\,2}=m^{2}\Rightarrow p^{0}=\pm E_{*}(\vec{p})\Rightarrow E_{*}(\vec{p})=\sqrt{\vec{p}^{\,2}+m^{2}}.$ (63)
@@
$v_{g}=\left|\frac{\partial E_{*}(\vec{p})}{\partial\vec{p}}\right|=\frac{|\vec{p}|}{\sqrt{\vec{p}^{\,2}+m^{2}}}<1,$ (64)
@@ Appendix A Supplementary Material
Therefore, we see that vectors in $\text{Im}\,Z$ and $\ker Z$ are eigenvectors of $S$. Now, let us show that $\text{Im}\,Z$ and $\ker Z$ cannot be trivial subspaces! For this to be true, we must have $\dim(\text{Im}\,Z)=\dim(\ker Z)=1$. Suppose initially that one of them is trivial and we will arrive at a contradiction.
Case 1: $\text{Im}\,Z=W$. Thus, $\forall\,w\in W$, we have
@@
Case 2: $\ker Z=W$. Thus, $\forall\,w\in W$, we have
@@
In general, $S_{W}=a\mathbf{1}_{W}$. Since $v=\omega(v)+\theta(v)$, it follows that
@@
$S(v)=a\mathbf{1}_{W}(\theta(v))=a\theta(v),$
@@
Since $\theta$ is symmetric and $S_{\mu\nu}$ is skew-symmetric, we obtain
$S_{\mu\nu}=a\theta_{\nu\mu}=S_{\nu\mu}=-S_{\mu\nu},$
which implies $S_{\mu\nu}=\varepsilon_{\mu\nu\beta}\,p^{\beta}=0$, and so $p=0$. But this is absurd, since we suppose $p\neq 0$. $\blacksquare$